PVFF Action Plan
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
View only
 
 
Still loading...
ABCEFGHIJKLMN
1
ID
Relates to ID
Date Raised
Parish
Assigned To (Organisation)
Assigned To (Person)
Original Due Date
Latest Due Date
Action
Last Updated
StatusStatus NotesFlood Forum Comments
2
11.11112/01/2016
All Parishes
PVFFPiers AllisonCR to arrange for a technical meeting between WBC and PA for how to share asset and map information18/05/2017In progressContacts put in place. But PA to follow with WBC contact (Stuart Powling)18/05/2017 CR Suggest close for WBC
3
3201/01/2014
Stanford Dingley
TWDetermine location of ingress to sewer network in Stanford Dingley.23/05/2017In progressDepth logger data has been reviewed but  not conclusive for points of ingresss. TW have confirmed they will now be trying smoke testing.
24/01/2017 work to line a sewer to be completed by 02/2017. However there remain problems in SD that require regular tanker visits and thus TW are carrying out further investigations via manhole lifting. 2 of 4 loggers are still working.

23/05/2017 TW to carry out an IAS survey when we have a wet winter. New Flygt pumps to be implement this summer in SD station.
Smoke test completed. As a result 4 pieces of work have now been raised by TW (3 in Frilsham and 1 in Stanford D) and now done.

18/5/2017 Meeting held in SD with BC, MH, KL and Andrew Hagger TW, Denise Kinsella TW and James Hoar TW.Maps of sewage pipework discussed and it was agreed that two flow monitors near Frilsham that were key to locating point of ingress were not working properly and possibly needed replacing. One data logger on Jennetts Hill to be moved up to Frilsham area to triangulate location of ingress. Some manholes still needing to be replaced on Jennetts Hill. An impermeable area survey was due to be done in Frilsham in approx 2-3 weeks time.  It was noted that it may still take some time to locate point of ingress and it may not be possible to do so until next winter and return of wetter weather. TW were keen to get to bottom of problem as SD was one of only two areas that had confounded them.
4
42.34215/09/2015BradfieldPVFFBradfield?Review operational arrangements for operating flow control structures at times of flooding at Bradfield mill and make provisions for back-up plan in case of operational difficulties during high flow events from flow control structures.24/01/2017In progressPVFF rep to investigate who the owner is (and contact details) and to document any suitable operating procedures.
24/01/2017 Owned by Bradfield College, operating being documented. We await to see if there are operating procedures (based on video of river levels)
23/05/2017 Paul to see Bradfield about this after the end of term.
5
454115/09/2015
All Parishes
PVFF01/10/2015Devise a village-centric account of causes, consequences and escalating risk levels. Mapping stage 212/01/2016On holdThis activity should be phase 2 of the Pang Valley flood risk mapping exercise
6
4815/09/2015West IlsleyWBCS Clark01/03/201631/03/2016West Ilsley to East Ilsley Road - investigate new road crossing at EA pump station.18/05/2017In progressTo be investigated in the 15/16 FY. No progress yet as of 12/01/2016.
20/09/2016 update: A pipe has been found. WBC to investigate opening up the pipe, liaising with the landowner on either side. This could not be done in FY16/17 as budget is committed.
24/01/2017 SC aims to look for funds in 17/18. PVFF Tech group will include this area in their slow sections.
23/05/2017 Pending suitable funds and staff availability.
No progress as at 18/05/2017.  Road crossing is the re-instatement of an existing culvert under Windmill Road. see 48.1
7
4915/09/2015West IlsleyWBCS Clark01/10/2015Investigate drainage and propose solutions at Hodcote House18/05/2017In progressInitial site visit and levels undertaken. Design on-going. Proposals submitted to owners Hodcote House. Feedback is "open ditch is dangerous for horses". SC (WBC) to discuss the proposal with Mick Channon - 2 options had been shown in proposals.
This is part of West Ilsley study – business case should be completed and submitted to the EA in June see ID 50
8
5015/09/2015West IlsleyWBCS Clark01/03/201631/03/2016Capital investigation into the upgrade of the surface water system through West Ilsley (Defra funded feasibility)18/05/2017In progressSite meeting held with consultant on 24th November. Feasibility has been completed. Level survey to be completed. To progress further SC must meet consultant on the cost of feasibility.
24/01/2017 This study is in progress and study should complete by end of 03/2017.
23/05/2017 Feasibility study in progress.
SC is still working on the business case.  Not 100% sure there is a viable solution but will carry on with the current proposal and submit the business case in June 2017. If we do have a viable scheme SC would look to do the detailed design next year (2018)and construct the year after (2019).
9
50.35012/01/2016West IlsleyWBC31/03/2016WBC to create a PAR for EA approval18/05/2017In ProgressSC about to start negotiating cost of PAR as at 10/3/16
24/01/2017 This study is in progress and study should complete by end of 03/2017.
PAR should be completed by end of Sept 2017
10
51.15112/01/2016West IlsleyTW01/01/2017TW to create a whole drainage strategy for West Ilsley07/09/2016In progressWest Ilsley Drainage strategy is within East Ilsley Drainage Strategy as the name is based on the location of the sewage treatment works. Stage 1 is on the TW website. Stage 2 due 1/9/2016

23/05/2017 Survey conducted in Compton but then it was dry. Not the best data. Awaiting wetter weather.
Now there won't be a Stage 2 report - there will be a combined Stage 2 and 3.  DK and TS will update us on progress at the next meeting
11
51.25112/01/2016West IlsleyTW01/02/2016TW to send out questionnaires to West Ilsley residents who were flooded. Need to confirm with PVFF reps to whom to send them. (Also East Ilsley)07/09/2016In progressTim Simms (Eight20)has had meeting with parish contacts for East Ilsley and Compton to determine where to send questionnaires. 90 will go to EI and 66 to Compton by 31/5/2016
20/09/2016: Questionnaires have been gathered. They are still to be reviewed. That information will inform the business case analysis. Note that elsewhere, such surveys have not helped the business case. Also carrying out CCTV and lift and look exercises.

23/05/2017 CCTV surveys scheduled, but need more groundwater for the work to succeed.18/5/2017
Questionnaires almost completed in E and W Ilsley. Compton scheduled for next round of surveys, start date awaited.
12
5215/09/2015West IlsleyWBCJ Winstanley01/10/2015Re-shape field to the rear of the Cottages24/01/2017On holdWBC to contact landowner. No progress as of 12/01/2016. Linked to SC (WBC) plan.
20/09/2016: no further progress. PAR study (see 50.3) will address this element has not progressed.
Awaiting a site meeting. No progress as at 7/9/2016
13
5515/09/2015ComptonTWLiam CrookOn-goingInvestigate Improving the management of surface water within former TWU abstraction facility on Ilsley Road to ease access for residents without any increase in flow downstream of this location.17/09/2016In progressDitches completed by WBC and TW.   24/01/2017 Flow monitors to be put in by TW.

23/05/2017 CCTV surveys scheduled, but need more groundwater for the work to succeed.
14
5815/09/2015
Hampstead Norreys
WBCS Clark01/03/2016Investigate a permanent overflow channel around Hampstead Norreys as provided during the flooding.18/05/2017In progress1. Overflow swale in field, 2. Blockage on bridge on B4009(underground channel to water meadow) – this wasn’t part of the WBC bid and there is a restriction below the B4009, 3. Sewage ingress was missed. WBC resubmitted a project mandate for the works - pending a response.
23/05/2017 Preliminary Design and Business Case 2017/18, Detailed design 2018/19, construction 2019/20
Hampstead Norreys – Preliminary Design and Business Case 2017/18, Detailed design 2018/19, construction 2019/20
15
5915/09/2015
Stanford Dingley
WBCS Clark01/03/2016Bid submitted to Defra (RFCC Grant Funding) for permanent bypass channel around Stanford Dingley. Sub-task - Complete investigation.18/05/2017In progressBid submitted through EA, awaiting confirmation of RFCC programme. Funding awarded for the investigation and feasibility in the 2015/16 financial year. UPDATE - The estimated cost of the study is higher than the grant allocation. Application for additional grant money submitted to the EA. It is Surface Water flooding. WBC have found money to conduct the study, design and create the PAR (aka the business case). PAR work has started with consultant working on this.
20/09/2016: when the consultants have a design and a model, SC to share with Stanford Dingley (Mark) by end of 03/2017
23/05/2017 PAR would allow 75 year protection. Business case stacks up.
.The scheme has been allocated 215k this year in the current EA Medium term programme. SC will begin the detailed design June 2017 . The current cost estimate for design and construction is 450k which includes the cost of lowering existing utilities underneath road. SC will need to discuss whether further funds will be available to take the scheme forward but the business case still stacks up.
16
63.26323/05/2017
All Parishes
PVFFPVFF to review the latest draft plan and comment back to WBC23/05/2017NewCR to sent to Kay to send out to others
17
6612/01/2016ComptonTW01/01/2017TW to create a whole drainage strategy for Compton07/09/2016In progressCompton Drainage Strategy Stage 1 on TW website
20/09/2016: TW will implement flow monitors in Compton to see if there is infiltration. Also planning to do CCTV work. Depth monitors in place, also in Hampstead Norreys.
24/01/2017 Update is the same for 20/09/2016 - current low water levels prevent data gathering.
TW to share trigger level for groundwater level when leaks might start into the sewer system. Installing rising main monitors East Ilsley pumping station and HN pumping station (budget is committed).

23/05/2017 CCTV surveys scheduled, but need more groundwater for the work to succeed.
The next step will be a combined Stage 2 and 3.  DK and TS will update on progress at the next PVFF meeting.
18
7220/09/2016
All Parishes
PVFFFlood mappersFlood Mappers should endeavour to maps as lines, TW sewers - where relevant (WI, EI, Compton, HN, SD)20/09/2016In progress
19
7520/09/2016TidmarshPVFFG DavidsonGeorge to agree operating procedures for Tidmarsh Mill (update for when owners are away)20/09/2016In progress23/05/2017 There is now an mitigation approach. George to update the procedures document and send back for publishing.
20
7724/01/2017
All Parishes
PVFFPAAdd to website the EA emergency contact details24/01/2017On hold
21
78.27823/01/2017
All Parishes
EAMWEA to share results of the CH2M modelling for the Pang and Sulham Brook in the Pangbourne and Tidmarsh areas.23/01/2017In progressMeeting to be arranged with Pangbourne project team re modelling CH2M
22
7924/01/2017
Pangbourne
WBCWBC to get a quotation for clearing gravel and debris from the downstream millstream at Pangbourne mill by end of 03/201718/05/2017On holdThe build up of slit downstream of the mill has been removed. A budget estimate to clear and remove silt immediately upstream of the mill would be about £5k, plus another £4k to replace the trash screen.
23/05/2017 It may be that further work is not needed. Initial clearing may have helped. Pangbourne PC to observe.
23
8024/01/2017
All Parishes
WBCWBC to investigate if the Jacobs report might change if the rain levels were increased by 30%18/05/2017On hold18/05/2017 No progress yet.
23/05/2017 Put on hold until National Goundwater study is published.
24
8124/01/2017
All Parishes
PVFFPVFF to map where temporary defences could/should be deployed.24/01/2017On hold
25
8224/01/2017
All Parishes
PVFFG DavidsonDraft a flood warden role document to circulate across the forum24/01/2017New
26
8424/01/2017TidmarshBBOWTBank strengthening work was done. The work was not done well and needs to be completed and to be fenced off. Kay to raise with them.18/05/2017In progressCL has checked the re-instated bank 20/5/2017 and happy that it is now re-vegetating and it also appears to be partly fenced off by BBOWT staff. However at the lasher weir the trash screen is still missing.  CL to follow up with BBOWT.  The logs on the bank also need to be secured with hoops of rebar.  GD and CL will do this.
27
8528/01/2017
Stanford Dingley
WBCJ Bowden/ P HenryWBC to organise maintenance of ditches in the water meadow next to The Bulll.18/05/2017In progressRiparian Owners were contacted by letter in 2016 suggesting that WBC organise the clearance on behalf of everyone with payment then being made to the Council. Some positive responses were received but we did not hear back from everyone.
Since then, site conditions initially prevented any work being done and currently due to other workload we do not have the capacity to progress this.
23/05/2017 Mark Hawkesworth to contact Riparian owners if J Bowden needs help.
28
85.18528/01/2017
Stanford Dingley
WBCJ Bowden/ P HenryWBC to seek contributions from a list of riparian owners towards the cost of this work18/05/2017In progressNot all riparian owners responded.  Site conditions now not suitable. see 85
29
87.18723/01/2017
All Parishes
PVFFPeter McGeehinPM to seek written confirmation on whether WBC, TW will commit funding to assist further research beyond the MSC (as EA (DB) confirmed that they would if the others did).23/01/2017NewMSC is due to finish on 23/08/2017. PM to receive the MSC thesis that can then be shared.
30
8823/05/2017
All Parishes
PVFFEach parish to check their own gully cleaning rota on the WBC website. If nothing cleaned then raise via the website. If then you get a bad response, raise to the complaints procedure. If you have frequently clogging gullies then consider raising for alternative solutions.23/05/2017New
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
60
61
62
63
64
65
66
67
68
69
70
71
72
73
74
75
76
77
78
79
80
81
82
83
84
85
86
87
88
89
90
91
92
93
94
95
96
97
98
99
100
Loading...
 
 
 
All Open Actions
All Actions
WBC Open Actions
TW Open Actions
EA Open Actions
PVFF Open Actions
Update Log
Data